Please simplify and show steps to the problem 20x^3y/4xy * -6xy/-x
Answer by Theo(13342)

see below the picture for further comments.

$$$

20x^3/4xy becomes 5x^2/y because 20/4 = 5 and x^3/x = x^2.

-6xy/-x becomes 6y because -x/-x = 1

5x^2*6 becomes 30x^2 becaose 5*6 = 30

the rest should be self explanatory.

write if you have any questions about this.